The moving-boundary method was applied to a 0.02000 mol/dm3 aqueous NaCl solution at 25°C using CdCl2 as the following solution. For a current held constant at 1.600 mA, Longsworth found that the boundary moved 10.00 cm in 3453 s in a tube of average cross-sectional area 0.1115 cm2. The conductivity of this NaCl solution at 25°C is 2.313 × 10−3 Ω−1 cm−1. Calculate u(Na+) and t(Na+) in this solution.
1
Expert's answer
2020-05-05T03:38:00-0400

Osmotic pressure = 4.26 atm


T = 27 + 273 = 300 K


C = 0.1


R = 0.821 L atm K^-1 mol^-1


π = i CRT


i = π / CRT = 4.26 / 0.1 x 0.821 x 300 = 0.173


i = Observed moles / Total moles = 1 + α / 1


i = 1 + α


0.173 = 1 + α


α = 1 - 0.173 = 0.827 = 82.7 %




Read more on Brainly.in - https://brainly.in/question/8932565#readmore